If we define x*y to mean 3x+2y, and x^y to mean 4x-y, then the value of (-2^3)*(3^-2) is
(A) -55 (B) -5 (C) 0 (D) 5 (E) 20

Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52778)

.

Follow to the standard rule: first make operations in parentheses.



Answer by math_tutor2020(3816)


x^y = 4x-y
(-2)^(3) = 4(-2)-3
(-2)^(3) = -8-3
(-2)^(3) = -11
Let A = (-2)^(3) = -11

x^y = 4x-y
(3)^(-2) = 4(3)-(-2)
(3)^(-2) = 12+2
(3)^(-2) = 14
Let B = (3)^(-2) = 14

x * y = 3x+2y
A * B = 3A+2B
A * B = 3(-11)+2(14)
A * B = -33+28
A * B = -5

Answer: (B) -5